# Non-polynomial divided difference and blossoming

### v1

@@ 2 Non-polynomial divided differences
that never vanishes for distinct $x_{1},x_{2}\in[a,b]$, Zürnacı and Dişibüyük give an explicit representation of non-polynomial B-spline functions for the spaces $\pi_{n}(\gamma_{1},\gamma_{2})$ which are a wide collection of spline spaces. These non-polynomial B-splines are constructed by using non-polynomial divided differences applied to a proper generalization of truncated-power function. In [11], starting with the interpolation problem, non-polynomial divided differences are defined recursively as a generalization of classical divided differences by the following definition.
@@ 3 Blossoming
    $p(x_{1},\ldots,(1-\alpha)u+\alpha w,\ldots x_{n})=(1-\alpha)p(x_{1},\ldots,u,\ldots x_{n})+\alpha p(x_{1},\ldots,w,\ldots x_{n})$
@@
    $p((x_{1},w_{1}),\ldots,\alpha(x_{i},w_{i})+\beta(x_{i},v_{i}),\ldots(x_{n},w_{n}))=\alpha p((x_{1},w_{1}),\ldots,(x_{i},w_{i}),\ldots(x_{n},w_{n}))$
    $+\beta p((x_{1},w_{1}),\ldots,(x_{i},v_{i}),\ldots(x_{n},w_{n}))$
@@
The extended blossom of order $k\in\mathbb{Z}$ of a function $F(x)$ is a function $f(x_{1},\ldots,x_{n+k}/v_{1},\ldots,x_{k})$ which satisfies the following axioms:
@@
    $f(x_{1},\ldots,x_{m}/v_{1},\ldots,x_{n})=f(x_{\sigma(1)},\ldots,x_{\sigma(m)}/v_{\tau(1)},\ldots,v_{\tau(n)})$
@@
    $f(x_{1},\ldots,(1-\alpha)u+\alpha v,\ldots x_{m}/v_{1},\ldots,x_{n})=(1-\alpha)f(x_{1},\ldots,u,\ldots x_{m}/v_{1},\ldots,x_{n})$
    $+\alpha f(x_{1},\ldots,v,\ldots x_{m}/v_{1},\ldots,x_{n}))$
@@
    $f(x_{1},\ldots,x_{m},w/v_{1},\ldots,x_{n},w)=f(x_{1},\ldots,x_{m}/v_{1},\ldots,x_{n})$
@@
    $g((x_{1},w_{1}),\ldots,(x_{n},w_{n}))=g((x_{\sigma(1)},w_{\sigma(1)}),\ldots,(x_{\sigma(n)},w_{\sigma(n)})$
@@
    $g((x_{1},w_{1}),\ldots,\alpha(x_{i},w_{i})+\beta(x_{i},v_{i}),\ldots(x_{n},w_{n}))=\alpha g((x_{1},w_{1}),\ldots,(x_{i},w_{i}),\ldots(x_{n},w_{n}))$
    $+\beta g((x_{1},w_{1}),\ldots,(x_{i},v_{i}),\ldots(x_{n},w_{n}))$
@@ 4 Non-polynomial divided difference and blossoming
$g_{e}((x_{1},w_{1}),\ldots,(x_{m},w_{m})/(u_{1},v_{1}),\ldots,(u_{n},v_{n}))=\sum(-1)^{\beta}g(x_{i_{1}},w_{i_{1}}),\ldots,(x_{i_{\alpha}},w_{i_{\alpha}}),(u_{j_{1}},v_{j_{1}}),\ldots,(u_{j_{\beta}},v_{j_{\beta}}),$ (4.3)
@@
It is necessary to verify that $g_{e}$ satisfies the axioms of the extended non-polynomial homogeneous blossom of order $k$. Using the axioms of the standard homogeneous blossom, it is clear that $g_{e}$ is bisymmetric and multiaffine in $(x,w)$. Assume, without loss of generality, that $(x_{1},w_{1})=(u_{1},v_{1})$. Then, by symmetry,
@@
$\bigg\{\left(D_{\gamma_{1},\gamma_{2}}d(a,x)\right)^{m}H(x)\bigg\}_{\gamma_{1},\gamma_{2}}\big[\varepsilon_{1},\ldots,\varepsilon_{n}\big]=h(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{m}/(u_{1},v_{1}),\ldots,(u_{n},v_{n})),$ (4.5)
where $(u_{i},v_{i})=(\gamma_{1}(\varepsilon_{i}),\gamma_{2}(\varepsilon_{i})),\;i=1,\ldots,n$ and $(\delta_{1},\delta_{2})=(D_{\gamma_{1},\gamma_{2}}(\gamma_{1}(a)),D_{\gamma_{1},\gamma_{2}}(\gamma_{2}(a)))$, $a\in\mathbb{Z}$.
@@
$\Psi(x)=(\delta_{1}\gamma_{2}(x)-\delta_{2}\gamma_{1}(x))^{m}.$
@@
$D_{\gamma_{1},\gamma_{2}}d(a,x)=D_{\gamma_{1},\gamma_{2}}[\gamma_{1}(a)\gamma_{2}(x)-\gamma_{2}(a)\gamma_{1}(x)]=\delta_{1}\gamma_{2}(x)-\delta_{2}\gamma_{1}(x),$
@@
$\binom{m}{j}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})=(d(x,a))^{j}\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}$ (4.6)
$\binom{k}{j}g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}/\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})=(d(x,a))^{j}\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}$ (4.7)
where $k=m-n\geq 0$, $(\delta_{1},\delta_{2})=(D_{\gamma_{1},\gamma_{2}}(\gamma_{1}(a)),D_{\gamma_{1},\gamma_{2}}(\gamma_{2}(a)))$, $a\in\mathbb{Z}$.
@@
$\Psi(\tau)=\prod_{i=1}^{m}\left(x_{i}\gamma_{2}(\tau)-w_{i}\gamma_{1}(\tau)\right).$
For the first $j$ elements of equation (4.8), taking $(x_{i},w_{i})=(\gamma_{1}(a),\gamma_{2}(a)),\quad i=1,\ldots,j$ yields that
$\Psi(\tau)=\left(d(a,\tau)\right)^{j}\prod_{i=j+1}^{m}\left(x_{i}\gamma_{2}(\tau)-w_{i}\gamma_{1}(\tau)\right).$ (4.9)
For all $(x_{i},w_{i})=(\gamma_{1}(x),\gamma_{2}(x)),\quad i=j+1,\ldots,m$, we get
$\Psi(\tau)=\left(d(a,\tau)\right)^{j}\left(d(x,\tau)\right)^{m-j},$
@@
$g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})$ $=\sum_{i}\frac{(-1)^{m-i}}{m!}D_{\gamma_{1},\gamma_{2}}^{i}\left((d(a,\tau))^{j}(d(x,\tau))^{m-j}\right)D_{\gamma_{1},\gamma_{2}}^{m-i}G(\tau).$ (4.10)
The generalized derivative satisfies the following formulas (see [11])
$D_{\gamma_{1},\gamma_{2}}^{k}(d(a,x))^{n}=\frac{n!}{(n-k)!}(d(a,x))^{n-k}$
$D_{\gamma_{1},\gamma_{2}}^{n}(f(x)g(x))=\sum_{i=0}^{n}\binom{n}{i}D_{\gamma_{1},\gamma_{2}}^{n-i}(f(x))D_{\gamma_{1},\gamma_{2}}^{i}(g(x)).$
Considering these formulas and setting $\tau=x$, equation (4.10) turns into
$g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})$ $=(d(a,x))^{j}(-1)^{j}\frac{(m-j)!}{m!}D_{\gamma_{1},\gamma_{2}}^{j}G(x)$
$=\frac{(m-j)!}{m!}(d(x,a))^{j}D_{\gamma_{1},\gamma_{2}}^{j}G(x).$
Hence,
$\binom{m}{j}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})$ $=\binom{m}{j}\frac{(m-j)!}{m!}(d(x,a))^{j}D_{\gamma_{1},\gamma_{2}}^{j}G(x)$
$=\frac{(d(x,a))^{j}}{j!}D_{\gamma_{1},\gamma_{2}}^{j}G(x).$
@@
$g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}$ $/\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})$
$=g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j}/)$
$=g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j})$
For $k=m-n\geq 0$, we can use the identity (4.6)
$\binom{m-n}{j}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j})=(d(x,a))^{j}\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}$
@@
$\binom{k}{j}g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}/\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})=(d(x,a))^{j}\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}.$
@@
$\left\{(d(a,x))^{n-1}G(x)\right\}_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]$
$=(d(a,x))^{n-1}\left\{\frac{(d-n+1)!}{d!}\right\}\sum g^{(n-1)}((u_{j_{1}},v_{j_{1}}),\ldots,(u_{j_{d-n+1}},v_{j_{d-n+1}})),$ (4.13)
@@
By Theorem 4.7,
$\left\{(d(a,x))^{n-1}G(x)\right\}_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]=g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{n-1}/(u_{1},v_{1}),\ldots,(u_{n},v_{n})),$
@@
$\left\{(d(a,x))^{n-1}G(x)\right\}_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]$
@@
$\frac{d!}{(d-n+1)!}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{n-1},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{d-n+1})=(d(x,a))^{n-1}D_{\gamma_{1},\gamma_{2}}^{n-1}G(x).$

### v2

@@ 2 Non-polynomial divided differences
that never vanishes for distinct $x_{1},x_{2}\in[a,b]$, Zürnacı and Dişibüyük give an explicit representation of non-polynomial B-spline functions for the spaces $\pi_{n}(\gamma_{1},\gamma_{2})$ which are a wide collection of spline spaces. These non-polynomial B-splines are constructed by using non-polynomial divided differences applied to a proper generalization of the truncated-power function. In [11], starting with the interpolation problem, non-polynomial divided differences are defined recursively as a generalization of classical divided differences by the following definition.
@@ 3 Blossoming
    $p(x_{1},\ldots,(1-\alpha)u+\alpha w,\ldots,x_{n})=(1-\alpha)p(x_{1},\ldots,u,\ldots,x_{n})+\alpha p(x_{1},\ldots,w,\ldots,x_{n})$
@@
    $p((x_{1},w_{1}),\ldots,\alpha(x_{i},w_{i})+\beta(x_{i},v_{i}),\ldots,(x_{n},w_{n}))=\alpha p((x_{1},w_{1}),\ldots,(x_{i},w_{i}),\ldots,(x_{n},w_{n}))$
    $+\beta p((x_{1},w_{1}),\ldots,(x_{i},v_{i}),\ldots,(x_{n},w_{n}))$
@@
The extended blossom of order $k\in\mathbb{Z}$ of a function $F(x)$ is a function $f(x_{1},\ldots,x_{n+k}/v_{1},\ldots,v_{k})$ which satisfies the following axioms:
@@
    $f(x_{1},\ldots,x_{m}/v_{1},\ldots,v_{n})=f(x_{\sigma(1)},\ldots,x_{\sigma(m)}/v_{\tau(1)},\ldots,v_{\tau(n)})$
@@
    $f(x_{1},\ldots,(1-\alpha)u+\alpha v,\ldots,x_{m}/v_{1},\ldots,v_{n})=(1-\alpha)f(x_{1},\ldots,u,\ldots,x_{m}/v_{1},\ldots,v_{n})$
    $+\alpha f(x_{1},\ldots,v,\ldots,x_{m}/v_{1},\ldots,v_{n})$
@@
    $f(x_{1},\ldots,x_{m},w/v_{1},\ldots,v_{n},w)=f(x_{1},\ldots,x_{m}/v_{1},\ldots,v_{n})$
@@
    $g((x_{1},w_{1}),\ldots,(x_{n},w_{n}))=g((x_{\sigma(1)},w_{\sigma(1)}),\ldots,(x_{\sigma(n)},w_{\sigma(n)}))$
@@
    $g((x_{1},w_{1}),\ldots,\alpha(x_{i},w_{i})+\beta(x_{i},v_{i}),\ldots,(x_{n},w_{n}))=\alpha g((x_{1},w_{1}),\ldots,(x_{i},w_{i}),\ldots,(x_{n},w_{n}))$
    $+\beta g((x_{1},w_{1}),\ldots,(x_{i},v_{i}),\ldots,(x_{n},w_{n}))$
@@ 4 Non-polynomial divided difference and blossoming
$g_{e}((x_{1},w_{1}),\ldots,(x_{m},w_{m})/(u_{1},v_{1}),\ldots,(u_{n},v_{n}))=\sum(-1)^{\beta}g(x_{i_{1}},w_{i_{1}}),\ldots,(x_{i_{\alpha}},w_{i_{\alpha}}),(u_{j_{1}},v_{j_{1}}),\ldots,(u_{j_{\beta}},v_{j_{\beta}})),$ (4.3)
@@
It is necessary to verify that $g_{e}$ satisfies the axioms of the extended non-polynomial homogeneous blossom of order $k$. Using the axioms of the standard homogeneous blossom, it is clear that $g_{e}$ is bisymmetric and multilinear in $(x,w)$. Assume, without loss of generality, that $(x_{1},w_{1})=(u_{1},v_{1})$. Then, by symmetry,
@@
$H(x)_{\gamma_{1},\gamma_{2}}\big[\varepsilon_{1},\ldots,\varepsilon_{n}\big]=(-1)^{m}h(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{m}/(u_{1},v_{1}),\ldots,(u_{n},v_{n})),$ (4.5)
where $(u_{i},v_{i})=(\gamma_{1}(\varepsilon_{i}),\gamma_{2}(\varepsilon_{i})),\;i=1,\ldots,n$ and $(\delta_{1},\delta_{2})=(D_{\gamma_{1},\gamma_{2}}(\gamma_{1}(a)),D_{\gamma_{1},\gamma_{2}}(\gamma_{2}(a)))$, $a\in\mathbb{R}$.
@@
$\Psi(x)=(D_{\gamma_{1},\gamma_{2}}d(x,a))^{m}=(-1)^{m}.$
@@
$D_{\gamma_{1},\gamma_{2}}d(x,a)=D_{\gamma_{1},\gamma_{2}}[\gamma_{2}(a)\gamma_{1}(x)-\gamma_{1}(a)\gamma_{2}(x)]=\delta_{2}\gamma_{1}(x)-\delta_{1}\gamma_{2}(x),$
@@
$\binom{m}{j}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})=\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!},$ (4.6)
$\binom{k}{j}g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}/\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})=\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!},$ (4.7)
where $k=m-n\geq 0$, $(\delta_{1},\delta_{2})=(D_{\gamma_{1},\gamma_{2}}(\gamma_{1}(a)),D_{\gamma_{1},\gamma_{2}}(\gamma_{2}(a)))$, $a\in\mathbb{R}$.
@@
$g((x_{1},w_{1}),\ldots,(x_{m},w_{m}))=\sum_{i}\frac{(-1)^{m-i}}{m!}D_{\gamma_{1},\gamma_{2}}^{i}\Psi(\tau)D_{\gamma_{1},\gamma_{2}}^{m-i}G(\tau),$ (4.8)
@@
$\Psi(\tau)=\prod_{r=1}^{m}\left(x_{r}\gamma_{2}(\tau)-w_{r}\gamma_{1}(\tau)\right).$
For the first $j$ elements of equation (4.8), taking $(x_{r},w_{r})=(\delta_{1},\delta_{2}),\quad r=1,\ldots,j$ yields that
$\Psi(\tau)=(-1)^{j}\prod_{r=j+1}^{m}\left(x_{r}\gamma_{2}(\tau)-w_{r}\gamma_{1}(\tau)\right).$ (4.9)
For all $(x_{r},w_{r})=(\gamma_{1}(x),\gamma_{2}(x)),\quad r=j+1,\ldots,m$, we get
$\Psi(\tau)=(-1)^{j}\left(d(x,\tau)\right)^{m-j}.$
@@
$g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})$ $=\sum_{i}\frac{(-1)^{m-i+j}}{m!}D_{\gamma_{1},\gamma_{2}}^{i}\left((d(x,\tau))^{m-j}\right)D_{\gamma_{1},\gamma_{2}}^{m-i}G(\tau).$ (4.10)
The generalized derivative satisfies the formula (see [11])
$D_{\gamma_{1},\gamma_{2}}^{k}(d(a,x))^{n}=\frac{n!}{(n-k)!}(d(a,x))^{n-k}.$
Then using this formula, equation (4.10) turns into
$g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})=$ $\sum_{i}\frac{(-1)^{m-i+j}}{m!}\frac{(m-j)!}{(m-j-i)!}(d(x,\tau))^{m-j-i}D_{\gamma_{1},\gamma_{2}}^{m-i}G(\tau).$
Now we set $\tau=x$. Since $d(x,x)=0$, only the term with i=mji=m-j remains in the sum. Therefore,
$g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j})=\frac{(m-j)!}{m!}D_{\gamma_{1},\gamma_{2}}^{j}G(x).$
Multiplying both sides by $\binom{m}{j}$ yields equation (4.6) since $\binom{m}{j}\frac{(m-j)!}{m!}=\frac{1}{j!}$.
@@
$g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}$ $/\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})$
$=g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j}/)$
$=g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j}).$
After cancellation, the expression reduces to the ordinary non-polynomial homogeneous blossom with mnm-n arguments. Applying equation (4.6) with $m$ replaced by mnm-n yields
$\binom{m-n}{j}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-n-j})=\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}.$
@@
$\binom{k}{j}g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{j},\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{m-j}/\underbrace{(\gamma_{1}(x),\gamma_{2}(x)),\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{n})=\frac{D_{\gamma_{1},\gamma_{2}}^{j}G(x)}{j!}.$
@@
$G(x)_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]=\left\{\frac{(d-n+1)!}{d!}\right\}\sum g^{(n-1)}((u_{j_{1}},v_{j_{1}}),\ldots,(u_{j_{d-n+1}},v_{j_{d-n+1}})),$ (4.13)
@@
The proof is modeled on the argument in [3]. By Theorem 4.7,
$G(x)_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]=(-1)^{n-1}g_{e}(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{n-1}/(u_{1},v_{1}),\ldots,(u_{n},v_{n})),$
@@
$G(x)_{\gamma_{1},\gamma_{2}}\left[\varepsilon_{1},\ldots,\varepsilon_{n}\right]$
@@
$\frac{d!}{(d-n+1)!}g(\underbrace{(\delta_{1},\delta_{2}),\ldots,(\delta_{1},\delta_{2})}_{n-1},\underbrace{(\gamma_{1}(x),\gamma_{2}(x))\ldots,(\gamma_{1}(x),\gamma_{2}(x))}_{d-n+1})=D_{\gamma_{1},\gamma_{2}}^{n-1}G(x).$
@@
The results obtained in this paper reduce to the classical polynomial case. Consider the space $\pi_{n}(\gamma_{1},\gamma_{2})$ with the choice
$\gamma_{1}(x)=x,\qquad\gamma_{2}(x)=1.$
In this case we obtain
$d(x_{1},x_{2})=\gamma_{1}(x_{1})\gamma_{2}(x_{2})-\gamma_{1}(x_{2})\gamma_{2}(x_{1})=x_{1}-x_{2}.$
Consequently the generalized derivative $D_{\gamma_{1},\gamma_{2}}$ reduces to the ordinary derivative and the non-polynomial divided difference becomes the classical polynomial divided difference. Also, the non-polynomial homogeneous blossom coincides with the standard blossom of a polynomial. Moreover, the parameter pairs appearing in the blossom take the form $(u_{i},v_{i})=(\varepsilon_{i},1)$. Since in the polynomial case the blossom depends only on the first coordinate, this identity reduces exactly to the classical relation between divided differences and blossoming obtained by Goldman [2]. Thus Theorem 4.10 may be viewed as a direct extension of Goldman’s polynomial formula to the non-polynomial framework considered in this paper.